OpenStudy (anonymous):

hi everybody i have question on increasing and decreasing function f(x)=x^1/2 it increasing on [0,&) why 0 is including and x^3-3x^2+1 it increasing on (-&,0].[2,&) why here also including since -x2 - 2x + 15 decreasing on (-1, +∞) and 1 is not including

OpenStudy (anonymous):

f(x)=x^1/2 well, at first if x=0, then y=0 ──> (0,0) if x=∞ (I think that's what you meant by saying &) then the y would be still between 1 and 0, but it would be approaching zero. Lets say x is negative,\[knowing~~that~~,~~A^{1/2}=\sqrt[2]{A^{1}~}=\sqrt{A}\] the n we can say that all negative values will be undefined, so they are not part of the function... yeah, it would be [0 if x=0 and ∞) because the x can be any large positive value.

OpenStudy (anonymous):

So, zero is included when x=0.

OpenStudy (anonymous):

(0)^1/2=0

OpenStudy (anonymous):

on x^2 not including why i have second exam in few days when i can say including and not including and what the meaning of undfine

OpenStudy (anonymous):

student_basil